# An Optimisation Framework for Unsupervised Environment Design

### v1

@@ 2.4 Choice of Score Function
One prevailing alternative to regret is the learnability of a level (Rutherford et al., 2024). Learnabilty was initially written in terms of the agent’s solve rate of a level, but can also be written as the variance of agent successes. Considering a policy’s trajectory distribution on a level to be $\pi_{x}(\lambda)$, we denote the policy’s return on a trajectory $\tau$ to be $R(\tau,\lambda)$. We thus rewrite learnability as:
$\text{Var}_{\tau\sim\pi_{x}(\lambda)}\bigg{[}R(\tau,\lambda)\bigg{]}$ (4)
[…]
@@ 4.1 Core Optimisation Problem
where $\alpha>0$ is a temperature coefficient. Our justification is twofold: first, as per Chung et al. (2024) the agent needs to train on several different levels at each iteration, requiring the adversary’s distribution to have greater entropy instead of collapsing to a single level whose score is largest. We also introduce an additional justification: entropy regularisation ensures that $f$ is strongly concave in $y$, guaranteeing best-iterate convergence (Theorem 5.1). In the case of a general score function, the optimisation problem can instead be written as:
$\min_{x\in\mathcal{X}}~{}f(x,y)=-y^{T}\boldsymbol{J}(\pi_{x},\Lambda)\,,~{}\max_{y\in\mathcal{Y}}~{}g(x,y)=y^{T}\boldsymbol{s}(\pi_{x},\Lambda)+\alpha\mathcal{H}(y)\,.$ (7)
Because UED conventionally uses a nonlinear neural network to parameterise its policy (and value function), we have an objective that is nonconvex in $x$ and concave in $y$. Lin et al. (2020) have shown that under certain assumptions, we can guarantee (best-iterate) convergence using two-timescale stochastic gradient descent-ascent, which assumes a separation of learning rates. Using such results, we propose NonConvex-Concave optimisation for UED (NCC) as a theoretically based method for optimisation in the UED setting.
@@ 4.2 Method for Optimisation
To perform gradient based optimisation for the adversary, we construct score vector $\mathbf{s}$ before each iteration of RL training to construct $y$’s gradient. For the adversary, we perform projected gradient ascent constrained to the probability simplex, and for $x$ we perform unconstrained gradient descent. The training loop is summarised in Algorithm 1 (illustrated in Figure 1), where $\mathcal{P}_{\mathcal{X}}(\cdot)$ represents the euclidean projection onto set $\mathcal{X}$. For learning rates $\eta_{y}\gg\eta_{x}$, and stochastic gradient estimators $\hat{F}$ and $\hat{G}$ defined in Equations 9 and 10, our update rule can be summarised as the following:
$x^{t+1}=x^{t}-\eta_{x}\cdot\hat{F},~{}~{}y^{t+1}=\mathcal{P}_{\mathcal{Y}}\big{(}y^{t}+\eta_{y}\cdot\hat{G}\big{)}\text{.}$ (8)
@@ 5 Convergence Results
In order to guarantee convergence in the zero-sum setting, we use two-timescale stochastic gradient descent-ascent (Lin et al., 2020), as given by Algorithm 1, to find an approximate solution to the optimisation problem defined by Equation (6). We first make the necessary assumptions and definitions, and then state the guarantees of our theoretical algorithm.
@@ 5.1 Preliminaries and Assumptions
We denote $F=\nabla_{x}f(x,y)$, $G=\nabla_{y}f(x,y)$, and $H=\nabla f=(F,G)$. Moreover, we let $N=\left\lvert\Lambda\right\rvert$ be the number of levels and $\lambda_{i}$ be the $i$-th level. We denote trajectories by $\tau^{t}=(o^{t},a^{t},r^{t})$, with $\pi_{x}(\tau^{t})\coloneqq\pi_{x}(a^{t}|o^{t})$ for short. We write $\Psi^{t}(\tau,\lambda)$ for the baseline used in policy gradients, corresponding either to the advantage function $A^{\pi_{x}}(\tau^{t},\lambda)$ or the return $R(\tau^{t},\lambda)$ from time $t$ onwards. Finally, let $M$ be the batch size used for each gradient update.
In order to prove convergence guarantees, we need to make some basic regularity assumptions on the UED and policy network architecture. The first is standard for returns to be finite and for regularity conditions to hold, while the second can be guaranteed by clipping, normalizing or otherwise bounding network weights, as discussed in Appendix A.
@@
The agent policy $\pi_{x}$ is a $\zeta$-greedy policy parameterized by an $L$-Lipschitz and $K$-smooth function. The adversary distribution is constrained to the $\xi$-truncated probability simplex, namely $\mathcal{Y}=\Delta_{\xi}(\Lambda)\coloneqq\{y\in\Delta(\Lambda)\mid y_{i}\geq\xi\ \forall i\}$.
@@ 5.2 Convergence Guarantees
Under Assumptions 1, LABEL: and 2, the $M=1$ estimator $\hat{H}=(\hat{F},\hat{G})$ defined by Equations 9 and 10 has $\sigma^{2}$-bounded variance with
$\sigma^{2}=4R_{*}^{2}\left(N+\frac{T^{2}L^{2}}{\zeta^{2}}\right)\,.$
Moreover, the corresponding objective function $f(x,y)=\operatorname{\mathbb{E}}_{\lambda}\left[s(\pi_{x},\lambda)\right]+\alpha\mathcal{H}(y)$ is $\alpha$-strongly concave in $y$ and $\sigma$-Lipschitz. Finally, $f$ is $\ell$-smooth with
@@
Under Assumptions 1, LABEL: and 2, let $\sigma,\ell$ be the constants defined in Proposition 1, $\alpha$ the entropy temperature, and $\Delta=\max_{y}f(x^{0},y)-\max_{y}f(x^{*},y)$ the objective distance between initial and optimal policies. For learning rates $\eta_{x}=\Theta(\alpha^{2}/\ell^{3})$ and $\eta_{y}=\Theta(1/\ell)$, and a batch size $M=\Theta(\max\{1,\sigma^{2}\ell/\alpha\epsilon^{2}\})$, Algorithm 1 finds an $\epsilon$-stationary policy $\pi_{x^{*}}$ such that $\left\lVert\nabla_{x}\max_{y}f(x^{*},y)\right\rVert<\epsilon$ in
$O\left(\frac{\Delta\ell^{3}}{\alpha^{2}\epsilon^{2}}+\frac{2\ell^{3}}{\alpha\epsilon^{4}}\right)$
iterations.
@@ 6.2 Heuristics and General-Sum UED
Given a set of trajectories $\mathcal{M}$ on level $\lambda$, we compute the level-wise empirical mean $\mu_{\lambda}$ = $\frac{1}{|\mathcal{M}|}\sum_{\tau\in\mathcal{M}}R(\tau,\lambda)$, overall mean $\mu=\frac{1}{|\Lambda|}\sum_{\lambda\in\Lambda}\mu_{\lambda}$, level-wise empirical variance $\sigma^{2}_{\lambda}=(\frac{1}{|\mathcal{M}|}\sum_{\tau\in\mathcal{M}}R(\tau,\lambda)^{2})-\mu_{\lambda}^{2}$, and overall variance $\sigma^{2}=(\frac{1}{|\Lambda|}\sum_{\lambda}\mu_{\lambda}^{2})-(\frac{1}{|\Lambda|}\sum_{\lambda}\mu_{\lambda})^{2}$. Using the Gaussian probability density function $\mathcal{N}(\cdot|\mu,\sigma^{2})$, we get the generalised learnability score function:
$s(\pi_{x},\lambda)=\sigma_{\lambda}\cdot\mathcal{N}(\mu_{\lambda}|\mu,\sigma^{2})\text{.}$ (11)
In Appendix B.4 we repeat the score function analysis of Rutherford et al. (2024), demonstrating the score function’s effectiveness on Minigrid.
@@ 7.1 Experimental Setup
We test our policy on benchmarks from Rutherford et al. (2024), although we decline to use JaxNav as the single-agent setting’s results are highly saturated, and the multi-agent setting introduces additional optimisation challenges. Thus, we report results on Minigrid (Chevalier-Boisvert et al., 2023) using the implementation from Coward et al. (2024) and XLand-Minigrid (Nikulin et al., 2023). We refer the reader to Rutherford et al. (2024) for more details on the individual environments. Additionally, we show that our method can obtain competitive performance on a more complex benchmark, Craftax (Matthews et al., 2024).
In all provided plots, we show our contributions in bold font, and suffix “NCC” with the particular score function used (see Rutherford et al. (2024) for a further discussion of such score functions). “Reg”, “Learn”, and “PVL” correspond to regret, learnability, and positive value loss, respectively. For Minigrid, we use a bar plot for easier interpretation, whereas for XLand-Minigrid and Craftax we plot evaluation results over the course of training to additionally compare sample efficiency.
Experiments were written in JAX (Bradbury et al., 2018) and we discuss experimental details in Appendix B.6. Results are averaged across 10 seeds, and standard error from the mean is displayed in the plots. All experiments use PPO (Schulman et al., 2017) as the RL algorithm of choice, although we remark that in theory our method could be used with other algorithms.
@@ 7.2 Results
Figure 2: Mean solve rates with 95% confidence intervals on Minigrid, a common UED testbed.
@@
We observe improved performance on the Minigrid holdout set with 60 walls. For readability, we use a bar plot with 95% confidence intervals following the recommendations of Agarwal et al. (2021), but using mean instead of interquartile mean for consistency. Additionally, we use Minigrid to highlight our method’s efficacy given a regret oracle, and thus do not test NCC-PVL. [^3] We did use PLR-MaxMC instead of PLR-Reg however, because we did not see a improvement in empirical performance. Despite the wisdom from Rutherford et al. (2024) that learnability improves learning, we hypothesise that due to the zero-sum nature of regret it may still lead to more robust policies in the end.
Refer to caption Refer to caption
Craftax XLand-Minigrid
@@
We use our new generalisation of learnability to outperform the highest-performing UED baseline from Matthews et al. (2024, PLR-MaxMC), although we find that domain randomisation to perform slightly better than PLR. We remark that we find performance is stronger in Craftax with a static buffer, and we highlight this to mention that in environments with a higher density of “good” levels, it may not be necessary to use a dynamic buffer. [^4] In Craftax, the DR distribution of levels is the same as the evaluation distribution. We attribute similar performance across generalised SFL and NCC with learnability to the algorithms’ shared emphasis on levels with high learnability. For additional implementation details for our Craftax experiments see Appendix B.2.
@@ 9 Conclusion
In this paper, we examine the connection between UED and optimisation theory, and thereby develop a new framework for optimisation within UED. This framework unlocks the area’s first convergence guarantees, and suggests a new method that produces more robust policies. Following prior work, we use our theoretical analysis to develop a practical algorithm that obtains strong results on robustness evaluations. Ultimately, we believe that our work provides a gateway to the creation of practical robust RL methods with guarantees under reasonable assumptions, which will become increasingly important as UED gains prevalence.
@@ Appendix A Proof of Proposition 1
Under Assumptions 1, LABEL: and 2, the estimator $\hat{H}=(\hat{F},\hat{G})$ defined by Equations 9 and 10, with $M=1$, has $\sigma^{2}$-bounded variance with
$\sigma^{2}=4R_{*}^{2}\left(N+\frac{T^{2}L^{2}}{\zeta^{2}}\right)\,.$
Moreover, the corresponding objective function $f(x,y)=\operatorname{\mathbb{E}}_{\lambda}\left[s(\pi_{x},\lambda)\right]+\alpha\mathcal{H}(y)$ is $\alpha$-strongly concave in $y$ and $\sigma$-Lipschitz. Finally, $f$ is $\ell$-smooth with
@@
Recall Equations 9 and 10 with $M=1$ from the main text:
@@
$\hat{G}(x,y)$ $=\hat{s}(\pi_{x},\Lambda)+\alpha\nabla_{y}\mathcal{H}(y)\,,$
@@
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{H}(z)-H(z)\right\rVert^{2}\right]$ $=\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}(z)-F(z)\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{s}(\pi_{x},\Lambda)-s(\pi_{x},\Lambda)\right\rVert^{2}\right]$
$\leq\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}(z)\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{s}(\pi_{x},\Lambda)\right\rVert^{2}\right],.$
For the second term, we easily obtain, for $M=1$,
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{s}(\pi_{x},\Lambda)\right\rVert^{2}\right]\leq\sum_{i}\operatorname{\mathbb{E}}\left[\hat{s}(\pi_{x},\lambda_{i})^{2}\right]\leq 4NR_{*}^{2}$
for both $s=-J$ and $s=\text{Reg}$. For the first term, we invoke Lipschitzness and $\zeta$-greediness of the policy $\pi$. For any trajectory $\tau$ and any level $\lambda$, we have $\left\lvert\sum_{t}\Psi_{t}(\tau,\lambda)\right\rvert\leq 2TR_{*}$ for both $\Psi^{t}(\tau,\lambda)=R^{t}(\tau,\lambda)$ and $\Psi^{t}(\tau,\lambda)=A^{\pi_{x}}(\tau^{t},\lambda)$, which combined with
@@
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{H}(z)-H(z)\right\rVert^{2}\right]$ $\leq 4NR_{*}^{2}+\frac{4T^{2}R_{*}^{2}L^{2}}{\zeta^{2}}=\sigma^{2}$
as required.
@@
(3) Lipschitzness of $f$. Follows directly from part (1) of the proof by applying Jensen’s inequality:
$\left\lVert\nabla f(z)\right\rVert^{2}=\left\lVert\operatorname{\mathbb{E}}\big{[}\hat{H}(z)\big{]}\right\rVert^{2}\leq\operatorname{\mathbb{E}}\left[\lVert{\hat{H}(z)^{2}\rVert}\right]\leq\sigma^{2}\,.$
@@
$\nabla_{x}^{2}J(\pi_{x},\lambda)=\operatorname{\mathbb{E}}_{\tau}\bigg{[}\sum_{t}R^{t}(\tau,\lambda)\Big{(}\nabla_{x}\log\pi_{x}(\tau^{t})\nabla_{x}\log p(\tau\mid\pi_{x})^{T}+\nabla_{x}^{2}\log\pi_{x}(\tau^{t})\Big{)}\bigg{]}$
@@
Putting everything together, we obtain

### v2

@@ 2.4 Choice of Score Function
One prevailing alternative to regret is the learnability of a level (Rutherford et al., 2024). Assuming a binary-outcome setting, where the return $R(\tau,\lambda)$ is in $\{0,1\}$ for all trajectories $\tau$ and levels $\lambda$, learnability is defined as $s(\pi_{x},\lambda)=p(1-p)$, where $p$ is the agent’s solve rate $p=\operatorname{\mathbb{E}}_{\tau\sim\pi_{x}(\lambda)}\left[R(\tau,\lambda)\right]$. For binary outcomes, this can be rewritten as the variance of returns as follows:
$s(\pi_{x},\lambda)=\text{Var}_{\tau\sim\pi_{x}(\lambda)}\bigg{[}R(\tau,\lambda)\bigg{]}\,.$ (4)
@@ 4.1 Core Optimisation Problem
where $\alpha>0$ is a temperature coefficient. Our justification is twofold: first, as per Chung et al. (2024) the agent needs to train on several different levels at each iteration, requiring the adversary’s distribution to have greater entropy instead of collapsing to a single level whose score is largest. Second, entropy regularisation ensures that $f$ is strongly concave in $y$, guaranteeing best-iterate convergence (Theorem 5.1). For general score functions, the optimisation problem is given by:
$\min_{x\in\mathcal{X}}~{}f(x,y)=-y^{T}\boldsymbol{J}(\pi_{x},\Lambda)\ \ ,\ \ \max_{y\in\mathcal{Y}}~{}g(x,y)=y^{T}\boldsymbol{s}(\pi_{x},\Lambda)+\alpha\mathcal{H}(y)\,.$ (7)
Because UED conventionally uses a nonlinear neural network to parameterise its policy (and value function), we have an objective that is nonconvex in $x$ and strongly concave in $y$. Lin et al. (2020) have shown that under certain assumptions, we can guarantee (best-iterate) convergence using two-timescale stochastic gradient descent-ascent, which assumes a separation of learning rates. Using these results, we propose NonConvex-Concave optimisation for UED (NCC) as a theoretically based method for optimisation in the UED setting.
@@ 4.2 Method for Optimisation
To perform gradient based optimisation for the adversary, we construct the score vector $\mathbf{s}$ before each iteration of RL training to construct $y$’s gradient. For the adversary, we perform projected gradient ascent constrained to the probability simplex, and for $x$ we perform unconstrained gradient descent. The training loop is summarised in Algorithm 1 (illustrated in Figure 1), where $\mathcal{P}_{\mathcal{X}}(\cdot)$ represents the Euclidean projection onto set $\mathcal{X}$. For learning rates $\eta_{y}\gg\eta_{x}$, and stochastic gradient estimators $\hat{F}$ and $\hat{G}$ defined in Equations 9 and 10, our update rule can be summarised as the following:
$x^{t+1}=x^{t}-\eta_{x}\cdot\hat{F},~{}~{}y^{t+1}=\mathcal{P}_{\mathcal{Y}}\big{(}y^{t}+\eta_{y}\cdot\hat{G}\big{)}\,.$ (8)
@@ 5 Convergence Results
In order to guarantee convergence in the zero-sum setting, we use two-timescale stochastic gradient descent-ascent (Lin et al., 2020), as given by Algorithm 1, to find an approximate solution to the optimisation problem defined by Equation (6). We first make the necessary assumptions and definitions, and then state our main theorem.
@@ 5.1 Preliminaries and Assumptions
We denote $F=\nabla_{x}f(x,y)$, $G=\nabla_{y}f(x,y)$, and $H=\nabla f=(F,G)$. Moreover, we let $N=\left\lvert\Lambda\right\rvert$ be the number of levels and $\lambda_{i}$ be the $i$-th level. We denote trajectories by $\tau^{t}=(o^{t},a^{t},r^{t})$, with $\pi_{x}(\tau^{t})\coloneqq\pi_{x}(a^{t}|o^{t})$ for short. We write $\Psi^{t}(\tau,\lambda)$ for the estimator used in policy gradients, typically taken to be the discounted Q-function $Q_{\pi_{x}}^{t}(\tau,\lambda)=\gamma^{t}Q_{\pi_{x}}(s^{t},a^{t},\lambda)$, advantage function $A_{\pi_{x}}^{t}(\tau,\lambda)=\gamma^{t}A_{\pi_{x}}(s^{t},a^{t},\lambda)$ or return $R^{t}(\tau,\lambda)=\sum_{h=t}^{t}\gamma^{h}r^{h}$ from time $t$ onwards. The discount ensures that the policy gradient is a true gradient, as clarified by Nota & Thomas (2020).
In order to prove convergence guarantees, we need to make some basic regularity assumptions on the UED and policy network architecture. The first is standard for returns to be finite and for regularity conditions to hold, while the second can be guaranteed by clipping, normalising or otherwise bounding network weights, as discussed in Appendix A.
@@
The agent policy $\pi_{x}$ is a $\zeta$-greedy policy parameterized by an $L$-Lipschitz and $K$-smooth function with parameters $x$. The adversary distribution is constrained to the $\xi$-truncated probability simplex, namely $\mathcal{Y}=\Delta_{\xi}(\Lambda)\coloneqq\{y\in\Delta(\Lambda)\mid y_{i}\geq\xi\ \forall i\}$.
@@ 5.2 Convergence Guarantees
Under Assumptions 1, LABEL: and 2, the estimator $\hat{H}=(\hat{F},\hat{G})$ defined in Equations 9 and 10 has $\sigma^{2}_{M}$-bounded variance, where
$\sigma^{2}_{M}=\frac{4R_{*}^{2}}{M}\left(N+\frac{T^{2}L^{2}}{\zeta^{2}}\right)\,.$
Moreover, the corresponding objective $f(x,y)=y^{T}\boldsymbol{s}(\pi_{x},\Lambda)+\alpha\mathcal{H}(y)$ defined in Equation 6 is $\alpha$-strongly concave in $y$, Lipschitz, and $\ell$-smooth, where
@@
Under Assumptions 1, LABEL: and 2, let $\ell$ and $\sigma\coloneqq\sigma_{1}$ be the constants defined in Proposition 1, $\alpha$ the entropy temperature, and $\Delta=\max_{y}f(x^{0},y)-\max_{y}f(x^{*},y)$ the objective distance between initial and optimal policies. For learning rates $\eta_{x}=\Theta(\alpha^{2}/\ell^{3})$ and $\eta_{y}=\Theta(1/\ell)$, and a batch size $M=\Theta(\max\{1,\sigma^{2}\ell/\alpha\epsilon^{2}\})$, Algorithm 1 finds an $\epsilon$-stationary policy $\pi_{x^{*}}$, such that $\left\lVert\nabla_{x}\max_{y}f(x^{*},y)\right\rVert<\epsilon$, in a number of iterations given by
$O\left(\frac{\Delta\ell^{3}}{\alpha^{2}\epsilon^{2}}+\frac{2\ell^{3}}{\alpha\epsilon^{4}}\right)\,.$
@@ 6.2 Heuristics and General-Sum UED
Given a set of $M$ trajectories $\{\tau_{i}\}_{i=1}^{M}$ on level $\lambda\in\Lambda$, we compute the level-wise empirical mean $\mu_{\lambda}$ = $\frac{1}{M}\sum_{i=1}^{M}R(\tau_{i},\lambda)$, overall mean $\mu=\frac{1}{N}\sum_{\lambda\in\Lambda}\mu_{\lambda}$, level-wise empirical variance $\sigma^{2}_{\lambda}=(\frac{1}{M}\sum_{i=1}^{M}R(\tau_{i},\lambda)^{2})-\mu_{\lambda}^{2}$, and overall variance $\sigma^{2}=(\frac{1}{N}\sum_{\lambda}\mu_{\lambda}^{2})-(\frac{1}{N}\sum_{\lambda}\mu_{\lambda})^{2}$. Using the Gaussian probability density function $\mathcal{N}(\cdot|\mu,\sigma^{2})$, we get the generalised learnability score:
$s(\pi_{x},\lambda)=\sigma_{\lambda}\cdot\mathcal{N}(\mu_{\lambda}|\mu,\sigma^{2})\,.$ (11)
In Appendix B.4 we repeat the score function analysis of Rutherford et al. (2024), demonstrating the score function’s effectiveness on Minigrid. Generalised learnability thus allows for the use of generalised SFL (“Gen-SFL” in Figure 3), which obtains superior performance on Craftax.
@@ 7.1 Experimental Setup
We test our method on benchmarks from Rutherford et al. (2024) and report results on Minigrid (Chevalier-Boisvert et al., 2023), using the implementation from Coward et al. (2024), and XLand-Minigrid (Nikulin et al., 2023). We omit JaxNav, as the single-agent setting’s results are highly saturated and the multi-agent setting introduces additional optimisation challenges. We refer the reader to Rutherford et al. (2024) for more details on the individual environments, noting only that Minigrid is the only environment with a regret oracle. Additionally, we show that our method can obtain competitive performance on a more complex benchmark, Craftax (Matthews et al., 2024).
In all provided plots, we show our contributions in bold font, and suffix “NCC” with the particular score function used (see Rutherford et al. (2024) for a further discussion of such score functions). “Reg”, “Learn”, and “PVL” correspond to regret, learnability, and positive value loss respectively. PVL is an approximation of regret when the latter is unknown, as in XLand-Minigrid and Craftax. For Minigrid, learning curves overlap significantly and we use a bar plot for visual clarity, whereas for XLand-Minigrid and Craftax we plot evaluation results over the course of training to additionally compare sample efficiency.
Experiments were written in JAX (Bradbury et al., 2018) and we discuss experimental details in Appendix B.6. Results are averaged across 10 seeds, and standard error from the mean is displayed in the plots. All experiments use PPO (Schulman et al., 2017) as the RL algorithm of choice.
@@ 7.2 Results
Figure 2: Mean solve rates with standard error bars on Minigrid, a common UED testbed.
@@
We observe improved performance on the Minigrid holdout set with 60 walls. For readability, we use a bar plot using recommendations from (Agarwal et al., 2021), but with mean and standard error for consistency. Additionally, we use Minigrid to highlight our method’s efficacy given a regret oracle, and thus do not test NCC-PVL. [^3] We did use PLR-MaxMC instead of PLR-Reg however, because we did not see a improvement in empirical performance. Despite the wisdom from Rutherford et al. (2024) that learnability improves learning, we hypothesise that due to the zero-sum nature of regret it may still lead to more robust policies in the end.
(a) XLand-Minigrid
(b) Craftax
@@
We use our new generalisation of learnability to outperform the highest-performing UED baseline from Matthews et al. (2024, PLR-MaxMC). We remark that we find performance is stronger in Craftax with a static buffer, and we highlight this to mention that in environments with a higher density of “good” levels, it may not be necessary to use a dynamic buffer. [^4] In Craftax, the DR distribution of levels is the same as the evaluation distribution. We attribute similar performance across generalised SFL and NCC with learnability to the algorithms’ shared emphasis on levels with high learnability. See Appendix B.2 for additional implementation details for Craftax experiments.
@@ 9 Conclusion
In this paper, we examine the connection between UED and optimisation theory, thereby developing a new framework for optimisation within UED. This framework unlocks the area’s first convergence guarantees, and suggests a new method that produces more robust policies. Following prior work, we use our theoretical analysis to develop a practical algorithm that obtains strong results on robustness evaluations. Ultimately, we believe that our work provides a gateway to the creation of practical robust RL methods with guarantees under reasonable assumptions, which will become increasingly important as UED gains prevalence.
@@ Appendix A Proof of Proposition 1
Under Assumptions 1, LABEL: and 2, the estimator $\hat{H}=(\hat{F},\hat{G})$ defined in Equations 9 and 10 has $\sigma^{2}_{M}$-bounded variance, where
$\sigma^{2}_{M}=\frac{4R_{*}^{2}}{M}\left(N+\frac{T^{2}L^{2}}{\zeta^{2}}\right)\,.$
Moreover, the corresponding objective $f(x,y)=y^{T}\boldsymbol{s}(\pi_{x},\Lambda)+\alpha\mathcal{H}(y)$ defined in Equation 6 is $\alpha$-strongly concave in $y$, Lipschitz, and $\ell$-smooth, where
@@
We first derive the bounds for $M=1$. Recall Equations 9 and 10 from the main text:
@@
$\hat{G}(x,y)$ $=\hat{\boldsymbol{s}}(\pi_{x},\Lambda)+\alpha\nabla_{y}\mathcal{H}(y)\,,$
@@
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{H}(z)-H(z)\right\rVert^{2}\right]$ $=\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}(z)-F(z)\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}(\pi_{x},\Lambda)-\boldsymbol{s}(\pi_{x},\Lambda)\right\rVert^{2}\right]$
$\leq\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}(z)\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}(\pi_{x},\Lambda)\right\rVert^{2}\right],.$
For the second term, we easily obtain
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}(\pi_{x},\Lambda)\right\rVert^{2}\right]\leq\sum_{i}\operatorname{\mathbb{E}}\left[\hat{s}(\pi_{x},\lambda_{i})^{2}\right]\leq 4NR_{*}^{2}$
for both $s=-J$ and $s=\text{Reg}$. For the first term, we invoke Lipschitzness and $\zeta$-greediness of the policy $\pi$. For any trajectory $\tau$ and any level $\lambda$, we have $\left\lvert\sum_{t}\Psi^{t}(\tau,\lambda)\right\rvert\leq 2TR_{*}$ for any choice of estimator $\Psi^{t}\in\{R^{t},Q_{\pi_{x}}^{t},A_{\pi_{x}}^{t}\}$, which combined with
@@
$\operatorname{\mathbb{E}}\left[\left\lVert\hat{H}(z)-H(z)\right\rVert^{2}\right]$ $\leq 4NR_{*}^{2}+\frac{4T^{2}R_{*}^{2}L^{2}}{\zeta^{2}}=\sigma^{2}_{1}$
as required. The variance for arbitrary $M$ follows immediately as $\sigma_{M}^{2}=\sigma_{1}^{2}/M$.
@@
(3) Lipschitzness of $f$. First note that
$\left\lVert\nabla_{y}\mathcal{H}\right\rVert^{2}=\sum_{i}(1+\log y_{i})^{2}\leq N(1+\log\xi)^{2}\,.$
We combine this with Jensen’s inequality and part (1) of the proof above to obtain $L$-Lipschitzness:
$\left\lVert\nabla f\right\rVert^{2}$ $=\left\lVert\operatorname{\mathbb{E}}\big{[}\hat{H}\big{]}\right\rVert^{2}\leq\operatorname{\mathbb{E}}\left[\left\lVert\hat{H}\right\rVert^{2}\right]$
$=\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}+\alpha\nabla_{y}\mathcal{H}\right\rVert^{2}\right]$
$=\operatorname{\mathbb{E}}\left[\left\lVert\hat{F}\right\rVert^{2}\right]+\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}\right\rVert^{2}\right]+\alpha^{2}\left\lVert\nabla_{y}\mathcal{H}\right\rVert^{2}+2\alpha\operatorname{\mathbb{E}}\left[\left\lVert\hat{\boldsymbol{s}}\right\rVert\right]\left\lVert\nabla_{y}\mathcal{H}\right\rVert$
$\leq\sigma_{1}^{2}+\alpha^{2}N(1+\log\xi)^{2}+4\alpha NR_{*}(1+\log\xi)\eqqcolon L\,.$
@@
$\nabla_{x}^{2}J(\pi_{x},\lambda)=\operatorname{\mathbb{E}}_{\tau}\bigg{[}\sum_{t}R^{t}(\tau,\lambda)\Big{(}\nabla_{x}\log\pi_{x}(\tau^{t})\nabla_{x}\log p(\tau\mid\pi_{x})^{T}+\nabla_{x}^{2}\log\pi_{x}(\tau^{t})\Big{)}\bigg{]}\,,$
@@
Putting everything together, recalling that $\left\lvert R^{t}(\tau,\lambda)\right\rvert\leq R_{*}$ for all $t$, we obtain